OpenStudy (anonymous):

Solve the following system of equations. x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 ANSWERS : (4, 4, –1) (–3, –1, –5) (–3, 5, –2) (2, –4, –3)

OpenStudy (anonymous):

y + 3z = -16 its (-3,-1,-5)

OpenStudy (mathmate):

Instead of solving the system, you can choose the right answer quickly by adding up the three equations to get 5x+3z=30 and then substitute each of the four answers in the above equation to find out if it works. (4, 4, –1) : 5*4-3*1=17 no! (–3, –1, –5) : 5*(-3)-5*3=-30 yes or maybe (–3, 5, –2) : 5*(-3)-5*2=-25 no! (2, –4, –3) : 5*2-3*3=1 no! So there is only one good answer, the second one. If you find more than one good answer, you'd have to substitute in each equation separately.

OpenStudy (anonymous):

THANKS (: